# Deep Reinforcement Learning for Advanced Longitudinal Control and Collision Avoidance in High-Risk Driving Scenarios

### v1

@@ Abstract
Existing Advanced Driver-Assistance Systems primarily focus on the vehicle directly ahead, often overlooking potential risks from following vehicles. This oversight can lead to ineffective handling of high-risk situations, such as high-speed, closely-spaced, multi-vehicle scenarios where emergency braking by one vehicle might trigger a pile-up collision. To overcome these limitations, this study introduces a novel deep reinforcement learning-based algorithm for longitudinal control and collision avoidance. This proposed algorithm effectively considers the behavior of both leading and following vehicles. Its implementation in simulated high-risk scenarios, which involve emergency braking in dense traffic where traditional systems typically fail, has demonstrated the algorithm’s ability to prevent potential pile-up collisions, including those involving heavy-duty vehicles.
@@ I Introduction
Advanced Driver-Assistance Systems (ADASs) are technologies designed to enhance traffic safety by monitoring the vehicle’s surrounding environment and automatically intervening to prevent potential crashes should the drivers fail to respond appropriately [1, 2]. ADAS emerged in the 1970s with the adoption of the anti-lock braking system.[3]. Over the years, the progression of technology has enabled ADAS to expand its functionalities, including automatic emergency braking (AEB) [4], stability control[5], blind-spot detection[6, 7], lane departure warnings[8], adaptive cruise control (ACC) [9], and traction control[10]. These developments have significantly improved traffic safety.
@@ III-A Markov decision process
0:  Current time step leading vehicle sensor detection result $data_{f}$ and $data_{r}$, Ego vehicle position $(x_{m},y_{m})$, Ego vehicle velocity $v_{m}$
@@
4:        Acquire the middle ego RL vehicle’s position $(x_{m},y_{m})$, velocity $v_{m}$ and the approximate gap distance $d_{fm}$ $d_{mr}$ between Leading/following vehicle and ego middle vehicle
@@ IV-A Baseline ADAS model
A baseline ADAS model combining AEB and ACC is developed to evaluate the proposed algorithm. The Time-To-Collision (TTC) is used as the AEB and ACC engagement factor. The vehicles’ velocities are kept at the same to reduce the perturbations [28]. If the TTC between the ego vehicle and the leading vehicle is less than $1.4s$, the ego vehicle will activate the emergency brake system. The algorithm of the baseline ADAS model is shown in Algorithm 2.
@@
5:        Predict leading vehicle velocity $v_{f}$ by Kalman Filter
6:        Predict leading vehicle new position $(x_{fnew},y_{fnew})=(x_{f},y_{f})+v_{f}*\Delta t$
@@ IV-C Design of scenarios
It is easy to train the RL algorithm in normal driving scenarios, but as mentioned before, the ADAS should be capable of handing all driving situations including high-risk ones. The study designed several scenarios aimed at causing the collision not only with the leading vehicle but also with the following vehicle to assess the algorithm’s capability to address the high-risk conditions that the current baseline ADAS algorithms cannot manage. The study also examines an edge case involving the following vehicle, which is heavy (compared to baseline light vehicle) and thus more prone to causing collisions with the ego vehicle. These scenarios include emergency brake scenarios (scenarios 1 and 2) and multiple vehicle following scenario (scenarios 3). A visual representation of these scenarios is provided in Fig.2.
@@ V-A DDPG training result
The DDPG agent was trained within the designed environment. Initially, the number of the training episodes was set to 10,000 to evaluate whether the neural network was converging or overfitting, and to determine if the RL agent had successfully learned a policy capable of avoiding collisions in high-risk scenarios. The robustness of the training process was assessed using 30 different random seeds. Due to the randomness, the convergence of the DDPG model slightly differed. The dark line and light shading in the Fig.3. represent the mean reward and its standard deviation to $1500^{th}$ episode, respectively. The DDPG model converged at approximately the $400^{th}$ episode, with a reward of $22500$ if the vehicle successfully avoided any collisions. 3.
@@ V-C1 Emergency brake scenarios
(b) Highway emergency brake scenario 1

### v2

@@ Abstract
Existing Advanced Driver-Assistance Systems primarily focus on the vehicle directly ahead, often overlooking potential risks from following vehicles. This oversight can lead to ineffective handling of high-risk situations, such as high-speed, closely-spaced, multi-vehicle scenarios where emergency braking by one vehicle might trigger a pile-up collision. To overcome these limitations, this study introduces a novel deep reinforcement learning-based algorithm for longitudinal control and collision avoidance. This proposed algorithm effectively considers the behavior of both leading and following vehicles. Its implementation in simulated high-risk scenarios, which involve emergency braking in dense traffic where traditional systems typically fail, has demonstrated the algorithm’s ability to prevent potential pile-up collisions, including those involving heavy-duty vehicles.x
@@ I Introduction
Advanced Driver-Assistance Systems (ADASs) are technologies designed to enhance traffic safety by monitoring the vehicle’s surrounding environment and automatically intervening to prevent potential crashes should the drivers fail to respond appropriately [1, 2]. ADAS emerged in the 1970s with the adoption of the anti-lock braking system [3]. Over the years, the progression of technology has enabled ADAS to expand its functionalities, including automatic emergency braking (AEB) [4], stability control [5], blind-spot detection[6, 7], lane departure warnings [8], adaptive cruise control (ACC) [9], and traction control [10]. These developments have significantly improved traffic safety.
@@ III-A Markov decision process
0:  Current time step preceding front vehicle sensor detection result $data_{f}$ and $data_{r}$, Ego vehicle position $(x_{m},y_{m})$, Ego vehicle velocity $v_{m}$
@@
4:        Acquire the middle RL vehicle’s position $(x_{m},y_{m})$, velocity $v_{m}$ and the approximate gap distance $d_{fm}$ $d_{mr}$ between front/rear vehicle and ego middle vehicle
@@ IV-A Baseline ADAS model
A baseline ADAS model combining AEB and ACC is developed to evaluate the proposed algorithm. The Time-To-Collision (TTC) is used as the AEB and ACC engagement factor. The vehicles’ velocities are kept at the same to reduce the perturbations [28]. Suppose the TTC between the ego vehicle and the leading vehicle is less than $1.4s$ (The Federal Highway Administration Advanced Driving Assistant System guide required) [29]. In that case, the ego vehicle will activate the emergency brake system. The algorithm of the baseline ADAS model is shown in Algorithm 2.
@@
5:        Predict preceding vehicle velocity $v_{f}$ by Kalman Filter
6:        Predict preceding vehicle new position $(x_{fnew},y_{fnew})=(x_{f},y_{f})+v_{f}*\Delta t$
@@ IV-C Design of scenarios
It is easy to train the RL algorithm in normal driving scenarios, but as mentioned before, the ADAS should be capable of handing all driving situations including high-risk ones. The study designed several scenarios aimed at causing the collision not only with the leading vehicle but also with the following vehicle to assess the algorithm’s capability to address the high-risk conditions that the current baseline ADAS algorithms cannot manage, generated by digital twin environments [30]. The study also examines an edge case involving the following vehicle, which is heavy (compared to baseline light vehicle) and thus more prone to causing collisions with the ego vehicle. These scenarios include emergency brake scenarios (scenarios 1 and 2) and multiple vehicle following scenario (scenarios 3). A visual representation of these scenarios is provided in Fig. 2.
@@ V-A DDPG training result
The DDPG agent was trained within the designed environment. Initially, the number of the training episodes was set to 10,000 to evaluate whether the neural network was converging or overfitting, and to determine if the RL agent had successfully learned a policy capable of avoiding collisions in high-risk scenarios. The robustness of the training process was assessed using 30 different random seeds. Due to the randomness, the convergence of the DDPG model slightly differed. The dark line and light shading in the Fig. 3. represent the mean reward and its standard deviation to $1500^{th}$ episode, respectively. The DDPG model converged at approximately the $400^{th}$ episode, with a reward of $22500$ if the vehicle successfully avoided any collisions as shown in Fig. 3.
@@ V-C1 Emergency brake scenarios
(b) Highway emergency brake scenario 2